### v2

# 00footnotetext: Received xxxx Month xxxx

Abstract
We present a measurement of the integrated luminosity of $e^{+}e^{-}$ collision data collected with the BESIII detector at the BEPCII collider at a center-of-mass energy of $E_{\rm cm}=3.773$ GeV. The integrated luminosities of the data sets taken from December 2021 to June 2022, from November 2022 to June 2023, and from October 2023 to February 2024 are determined to be $4.995\pm 0.019$ fb-1, $8.157\pm 0.031$ fb-1, and $4.191\pm 0.016$ fb-1, respectively, by analyzing large angle Bhabha scattering events. The uncertainties are dominated by systematic effects and the statistical uncertainties are negligible. Our results provide essential input for future analyses and precision measurements.
keywords:
Bhabha scattering events, integrated luminosity, cross section
M. Ablikim1, M. N. Achasov4,c, P. Adlarson75, O. Afedulidis3, X. C. Ai80, R. Aliberti35, A. Amoroso74A,74C, Q. An71,58,a, Y. Bai57, O. Bakina36, I. Balossino29A, Y. Ban46,h, H.-R. Bao63, V. Batozskaya1,44, K. Begzsuren32, N. Berger35, M. Berlowski44, M. Bertani28A, D. Bettoni29A, F. Bianchi74A,74C, E. Bianco74A,74C, A. Bortone74A,74C, I. Boyko36, R. A. Briere5, A. Brueggemann68, H. Cai76, X. Cai1,58, A. Calcaterra28A, G. F. Cao1,63, N. Cao1,63, S. A. Cetin62A, J. F. Chang1,58, G. R. Che43, G. Chelkov36,b, C. Chen43, C. H. Chen9, Chao Chen55, G. Chen1, H. S. Chen1,63, H. Y. Chen20, M. L. Chen1,58,63, S. J. Chen42, S. L. Chen45, S. M. Chen61, T. Chen1,63, X. R. Chen31,63, X. T. Chen1,63, Y. B. Chen1,58, Y. Q. Chen34, Z. J. Chen25,i, Z. Y. Chen1,63, S. K. Choi10A, G. Cibinetto29A, F. Cossio74C, J. J. Cui50, H. L. Dai1,58, J. P. Dai78, A. Dbeyssi18, R.  E. de Boer3, D. Dedovich36, C. Q. Deng72, Z. Y. Deng1, A. Denig35, I. Denysenko36, M. Destefanis74A,74C, F. De Mori74A,74C, B. Ding66,1, X. X. Ding46,h, Y. Ding34, Y. Ding40, J. Dong1,58, L. Y. Dong1,63, M. Y. Dong1,58,63, X. Dong76, M. C. Du1, S. X. Du80, Y. Y. Duan55, Z. H. Duan42, P. Egorov36,b, Y. H. Fan45, J. Fang59, J. Fang1,58, S. S. Fang1,63, W. X. Fang1, Y. Fang1, Y. Q. Fang1,58, R. Farinelli29A, L. Fava74B,74C, F. Feldbauer3, G. Felici28A, C. Q. Feng71,58, J. H. Feng59, Y. T. Feng71,58, M. Fritsch3, C. D. Fu1, J. L. Fu63, Y. W. Fu1,63, H. Gao63, X. B. Gao41, Y. N. Gao46,h, Yang Gao71,58, S. Garbolino74C, I. Garzia29A,29B, L. Ge80, P. T. Ge76, Z. W. Ge42, C. Geng59, E. M. Gersabeck67, A. Gilman69, K. Goetzen13, L. Gong40, W. X. Gong1,58, W. Gradl35, S. Gramigna29A,29B, M. Greco74A,74C, M. H. Gu1,58, Y. T. Gu15, C. Y. Guan1,63, A. Q. Guo31,63, L. B. Guo41, M. J. Guo50, R. P. Guo49, Y. P. Guo12,g, A. Guskov36,b, J. Gutierrez27, K. L. Han63, T. T. Han1, F. Hanisch3, X. Q. Hao19, F. A. Harris65, K. K. He55, K. L. He1,63, F. H. Heinsius3, C. H. Heinz35, Y. K. Heng1,58,63, C. Herold60, T. Holtmann3, P. C. Hong34, G. Y. Hou1,63, X. T. Hou1,63, Y. R. Hou63, Z. L. Hou1, B. Y. Hu59, H. M. Hu1,63, J. F. Hu56,j, S. L. Hu12,g, T. Hu1,58,63, Y. Hu1, G. S. Huang71,58, K. X. Huang59, L. Q. Huang31,63, X. T. Huang50, Y. P. Huang1, Y. S. Huang59, T. Hussain73, F. Hölzken3, N. Hüsken35, N. in der Wiesche68, J. Jackson27, S. Janchiv32, J. H. Jeong10A, Q. Ji1, Q. P. Ji19, W. Ji1,63, X. B. Ji1,63, X. L. Ji1,58, Y. Y. Ji50, X. Q. Jia50, Z. K. Jia71,58, D. Jiang1,63, H. B. Jiang76, P. C. Jiang46,h, S. S. Jiang39, T. J. Jiang16, X. S. Jiang1,58,63, Y. Jiang63, J. B. Jiao50, J. K. Jiao34, Z. Jiao23, S. Jin42, Y. Jin66, M. Q. Jing1,63, X. M. Jing63, T. Johansson75, S. Kabana33, N. Kalantar-Nayestanaki64, X. L. Kang9, X. S. Kang40, M. Kavatsyuk64, B. C. Ke80, V. Khachatryan27, A. Khoukaz68, R. Kiuchi1, O. B. Kolcu62A, B. Kopf3, M. Kuessner3, X. Kui1,63, N.  Kumar26, A. Kupsc44,75, W. Kühn37, J. J. Lane67, L. Lavezzi74A,74C, T. T. Lei71,58, Z. H. Lei71,58, M. Lellmann35, T. Lenz35, C. Li47, C. Li43, C. H. Li39, Cheng Li71,58, D. M. Li80, F. Li1,58, G. Li1, H. B. Li1,63, H. J. Li19, H. N. Li56,j, Hui Li43, J. R. Li61, J. S. Li59, K. Li1, L. J. Li1,63, L. K. Li1, Lei Li48, M. H. Li43, P. R. Li38,k,l, Q. M. Li1,63, Q. X. Li50, R. Li17,31, S. X. Li12, T.  Li50, W. D. Li1,63, W. G. Li1,a, X. Li1,63, X. H. Li71,58, X. L. Li50, X. Y. Li1,63, X. Z. Li59, Y. G. Li46,h, Z. J. Li59, Z. Y. Li78, C. Liang42, H. Liang1,63, H. Liang71,58, Y. F. Liang54, Y. T. Liang31,63, G. R. Liao14, Y. P. Liao1,63, J. Libby26, A.  Limphirat60, C. C. Lin55, D. X. Lin31,63, T. Lin1, B. J. Liu1, B. X. Liu76, C. Liu34, C. X. Liu1, F. Liu1, F. H. Liu53, Feng Liu6, G. M. Liu56,j, H. Liu38,k,l, H. B. Liu15, H. H. Liu1, H. M. Liu1,63, Huihui Liu21, J. B. Liu71,58, J. Y. Liu1,63, K. Liu38,k,l, K. Y. Liu40, Ke Liu22, L. Liu71,58, L. C. Liu43, Lu Liu43, M. H. Liu12,g, P. L. Liu1, Q. Liu63, S. B. Liu71,58, T. Liu12,g, W. K. Liu43, W. M. Liu71,58, X. Liu38,k,l, X. Liu39, Y. Liu80, Y. Liu38,k,l, Y. B. Liu43, Z. A. Liu1,58,63, Z. D. Liu9, Z. Q. Liu50, X. C. Lou1,58,63, F. X. Lu59, H. J. Lu23, J. G. Lu1,58, X. L. Lu1, Y. Lu7, Y. P. Lu1,58, Z. H. Lu1,63, C. L. Luo41, J. R. Luo59, M. X. Luo79, T. Luo12,g, X. L. Luo1,58, X. R. Lyu63, Y. F. Lyu43, F. C. Ma40, H. Ma78, H. L. Ma1, J. L. Ma1,63, L. L. Ma50, M. M. Ma1,63, Q. M. Ma1, R. Q. Ma1,63, T. Ma71,58, X. T. Ma1,63, X. Y. Ma1,58, Y. Ma46,h, Y. M. Ma31, F. E. Maas18, M. Maggiora74A,74C, S. Malde69, Y. J. Mao46,h, Z. P. Mao1, S. Marcello74A,74C, Z. X. Meng66, J. G. Messchendorp13,64, G. Mezzadri29A, H. Miao1,63, T. J. Min42, R. E. Mitchell27, X. H. Mo1,58,63, B. Moses27, N. Yu. Muchnoi4,c, J. Muskalla35, Y. Nefedov36, F. Nerling18,e, L. S. Nie20, I. B. Nikolaev4,c, Z. Ning1,58, S. Nisar11,m, Q. L. Niu38,k,l, W. D. Niu55, Y. Niu 50, S. L. Olsen63, Q. Ouyang1,58,63, S. Pacetti28B,28C, X. Pan55, Y. Pan57, A.  Pathak34, Y. P. Pei71,58, M. Pelizaeus3, H. P. Peng71,58, Y. Y. Peng38,k,l, K. Peters13,e, J. L. Ping41, R. G. Ping1,63, S. Plura35, V. Prasad33, F. Z. Qi1, H. Qi71,58, H. R. Qi61, M. Qi42, T. Y. Qi12,g, S. Qian1,58, W. B. Qian63, C. F. Qiao63, X. K. Qiao80, J. J. Qin72, L. Q. Qin14, L. Y. Qin71,58, X. P. Qin12,g, X. S. Qin50, Z. H. Qin1,58, J. F. Qiu1, Z. H. Qu72, C. F. Redmer35, K. J. Ren39, A. Rivetti74C, M. Rolo74C, G. Rong1,63, Ch. Rosner18, S. N. Ruan43, N. Salone44, A. Sarantsev36,d, Y. Schelhaas35, K. Schoenning75, M. Scodeggio29A, K. Y. Shan12,g, W. Shan24, X. Y. Shan71,58, Z. J. Shang38,k,l, J. F. Shangguan16, L. G. Shao1,63, M. Shao71,58, C. P. Shen12,g, H. F. Shen1,8, W. H. Shen63, X. Y. Shen1,63, B. A. Shi63, H. Shi71,58, H. C. Shi71,58, J. L. Shi12,g, J. Y. Shi1, Q. Q. Shi55, S. Y. Shi72, X. Shi1,58, J. J. Song19, T. Z. Song59, W. M. Song34,1, Y.  J. Song12,g, Y. X. Song46,h,n, S. Sosio74A,74C, S. Spataro74A,74C, F. Stieler35, Y. J. Su63, G. B. Sun76, G. X. Sun1, H. Sun63, H. K. Sun1, J. F. Sun19, K. Sun61, L. Sun76, S. S. Sun1,63, T. Sun51,f, W. Y. Sun34, Y. Sun9, Y. J. Sun71,58, Y. Z. Sun1, Z. Q. Sun1,63, Z. T. Sun50, C. J. Tang54, G. Y. Tang1, J. Tang59, M. Tang71,58, Y. A. Tang76, L. Y. Tao72, Q. T. Tao25,i, M. Tat69, J. X. Teng71,58, V. Thoren75, W. H. Tian59, Y. Tian31,63, Z. F. Tian76, I. Uman62B, Y. Wan55, S. J. Wang 50, B. Wang1, B. L. Wang63, Bo Wang71,58, D. Y. Wang46,h, F. Wang72, H. J. Wang38,k,l, J. J. Wang76, J. P. Wang 50, K. Wang1,58, L. L. Wang1, M. Wang50, N. Y. Wang63, S. Wang12,g, S. Wang38,k,l, T.  Wang12,g, T. J. Wang43, W. Wang59, W.  Wang72, W. P. Wang35,71,o, X. Wang46,h, X. F. Wang38,k,l, X. J. Wang39, X. L. Wang12,g, X. N. Wang1, Y. Wang61, Y. D. Wang45, Y. F. Wang1,58,63, Y. L. Wang19, Y. N. Wang45, Y. Q. Wang1, Yaqian Wang17, Yi Wang61, Z. Wang1,58, Z. L.  Wang72, Z. Y. Wang1,63, Ziyi Wang63, D. H. Wei14, F. Weidner68, S. P. Wen1, Y. R. Wen39, U. Wiedner3, G. Wilkinson69, M. Wolke75, L. Wollenberg3, C. Wu39, J. F. Wu1,8, L. H. Wu1, L. J. Wu1,63, X. Wu12,g, X. H. Wu34, Y. Wu71,58, Y. H. Wu55, Y. J. Wu31, Z. Wu1,58, L. Xia71,58, X. M. Xian39, B. H. Xiang1,63, T. Xiang46,h, D. Xiao38,k,l, G. Y. Xiao42, S. Y. Xiao1, Y.  L. Xiao12,g, Z. J. Xiao41, C. Xie42, X. H. Xie46,h, Y. Xie50, Y. G. Xie1,58, Y. H. Xie6, Z. P. Xie71,58, T. Y. Xing1,63, C. F. Xu1,63, C. J. Xu59, G. F. Xu1, H. Y. Xu66,2,p, M. Xu71,58, Q. J. Xu16, Q. N. Xu30, W. Xu1, W. L. Xu66, X. P. Xu55, Y. C. Xu77, Z. S. Xu63, F. Yan12,g, L. Yan12,g, W. B. Yan71,58, W. C. Yan80, X. Q. Yan1, H. J. Yang51,f, H. L. Yang34, H. X. Yang1, T. Yang1, Y. Yang12,g, Y. F. Yang1,63, Y. F. Yang43, Y. X. Yang1,63, Z. W. Yang38,k,l, Z. P. Yao50, M. Ye1,58, M. H. Ye8, J. H. Yin1, Z. Y. You59, B. X. Yu1,58,63, C. X. Yu43, G. Yu1,63, J. S. Yu25,i, T. Yu72, X. D. Yu46,h, Y. C. Yu80, C. Z. Yuan1,63, J. Yuan34, J. Yuan45, L. Yuan2, S. C. Yuan1,63, Y. Yuan1,63, Z. Y. Yuan59, C. X. Yue39, A. A. Zafar73, F. R. Zeng50, S. H.  Zeng72, X. Zeng12,g, Y. Zeng25,i, Y. J. Zeng59, Y. J. Zeng1,63, X. Y. Zhai34, Y. C. Zhai50, Y. H. Zhan59, A. Q. Zhang1,63, B. L. Zhang1,63, B. X. Zhang1, D. H. Zhang43, G. Y. Zhang19, H. Zhang80, H. Zhang71,58, H. C. Zhang1,58,63, H. H. Zhang34, H. H. Zhang59, H. Q. Zhang1,58,63, H. R. Zhang71,58, H. Y. Zhang1,58, J. Zhang80, J. Zhang59, J. J. Zhang52, J. L. Zhang20, J. Q. Zhang41, J. S. Zhang12,g, J. W. Zhang1,58,63, J. X. Zhang38,k,l, J. Y. Zhang1, J. Z. Zhang1,63, Jianyu Zhang63, L. M. Zhang61, Lei Zhang42, P. Zhang1,63, Q. Y. Zhang34, R. Y. Zhang38,k,l, S. H. Zhang1,63, Shulei Zhang25,i, X. D. Zhang45, X. M. Zhang1, X. Y. Zhang50, Y.  Zhang72, Y. Zhang1, Y.  T. Zhang80, Y. H. Zhang1,58, Y. M. Zhang39, Yan Zhang71,58, Z. D. Zhang1, Z. H. Zhang1, Z. L. Zhang34, Z. Y. Zhang76, Z. Y. Zhang43, Z. Z.  Zhang45, G. Zhao1, J. Y. Zhao1,63, J. Z. Zhao1,58, L. Zhao1, Lei Zhao71,58, M. G. Zhao43, N. Zhao78, R. P. Zhao63, S. J. Zhao80, Y. B. Zhao1,58, Y. X. Zhao31,63, Z. G. Zhao71,58, A. Zhemchugov36,b, B. Zheng72, B. M. Zheng34, J. P. Zheng1,58, W. J. Zheng1,63, Y. H. Zheng63, B. Zhong41, X. Zhong59, H.  Zhou50, J. Y. Zhou34, L. P. Zhou1,63, S.  Zhou6, X. Zhou76, X. K. Zhou6, X. R. Zhou71,58, X. Y. Zhou39, Y. Z. Zhou12,g, J. Zhu43, K. Zhu1, K. J. Zhu1,58,63, K. S. Zhu12,g, L. Zhu34, L. X. Zhu63, S. H. Zhu70, T. J. Zhu12,g, W. D. Zhu41, Y. C. Zhu71,58, Z. A. Zhu1,63, J. H. Zou1, J. Zu71,58
(BESIII Collaboration)
1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
17 Hebei University, Baoding 071002, People’s Republic of China
18 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
19 Henan Normal University, Xinxiang 453007, People’s Republic of China
20 Henan University, Kaifeng 475004, People’s Republic of China
21 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
22 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
23 Huangshan College, Huangshan 245000, People’s Republic of China
24 Hunan Normal University, Changsha 410081, People’s Republic of China
25 Hunan University, Changsha 410082, People’s Republic of China
26 Indian Institute of Technology Madras, Chennai 600036, India
27 Indiana University, Bloomington, Indiana 47405, USA
28 INFN Laboratori Nazionali di Frascati , (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
29 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
30 Inner Mongolia University, Hohhot 010021, People’s Republic of China
31 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
32 Institute of Physics and Technology, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
33 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
34 Jilin University, Changchun 130012, People’s Republic of China
35 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
36 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
37 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
38 Lanzhou University, Lanzhou 730000, People’s Republic of China
39 Liaoning Normal University, Dalian 116029, People’s Republic of China
40 Liaoning University, Shenyang 110036, People’s Republic of China
41 Nanjing Normal University, Nanjing 210023, People’s Republic of China
42 Nanjing University, Nanjing 210093, People’s Republic of China
43 Nankai University, Tianjin 300071, People’s Republic of China
44 National Centre for Nuclear Research, Warsaw 02-093, Poland
45 North China Electric Power University, Beijing 102206, People’s Republic of China
46 Peking University, Beijing 100871, People’s Republic of China
47 Qufu Normal University, Qufu 273165, People’s Republic of China
48 Renmin University of China, Beijing 100872, People’s Republic of China
49 Shandong Normal University, Jinan 250014, People’s Republic of China
50 Shandong University, Jinan 250100, People’s Republic of China
51 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
52 Shanxi Normal University, Linfen 041004, People’s Republic of China
53 Shanxi University, Taiyuan 030006, People’s Republic of China
54 Sichuan University, Chengdu 610064, People’s Republic of China
55 Soochow University, Suzhou 215006, People’s Republic of China
56 South China Normal University, Guangzhou 510006, People’s Republic of China
57 Southeast University, Nanjing 211100, People’s Republic of China
58 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
59 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
60 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
61 Tsinghua University, Beijing 100084, People’s Republic of China
62 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
63 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
64 University of Groningen, NL-9747 AA Groningen, The Netherlands
65 University of Hawaii, Honolulu, Hawaii 96822, USA
66 University of Jinan, Jinan 250022, People’s Republic of China
67 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
68 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
69 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
70 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
71 University of Science and Technology of China, Hefei 230026, People’s Republic of China
72 University of South China, Hengyang 421001, People’s Republic of China
73 University of the Punjab, Lahore-54590, Pakistan
74 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
75 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
76 Wuhan University, Wuhan 430072, People’s Republic of China
77 Yantai University, Yantai 264005, People’s Republic of China
78 Yunnan University, Kunming 650500, People’s Republic of China
79 Zhejiang University, Hangzhou 310027, People’s Republic of China
80 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
p Also at School of Physics, Beihang University, Beijing 100191 , China
1 Introduction
Luminosity plays a crucial role in quantifying the size of a dataset and is a fundamental parameter for measuring various physics processes, particularly cross sections. The number of events for the process $e^{+}e^{-}\to X$ ($X$ denotes any possible final state) in $e^{+}e^{-}$ collision data can be expressed as
$N_{e^{+}e^{-}\to X}=\mathcal{L}\times\sigma_{e^{+}e^{-}\to X}(E_{\rm cm}),$ (1)
where $\mathcal{L}$ denotes the integrated luminosity of the data set, $\sigma_{e^{+}e^{-}\to X}$ is the cross section for the process $e^{+}e^{-}\to X$, and $E_{\rm cm}$ is the center-of-mass energy. In principle, any process with a known cross section can be used to determine luminosity. However, Quantum Electrodynamics (QED) processes are advantageous due to their high production rates, simple final state topologies, and cross sections that are known with high theoretical precision.
Datasets at $E_{\rm cm}=3.773$ GeV were collected by the BESIII detector at the BEPCII collider from December 2021 to June 2022 (DATA I), from November 2022 to June 2023 (DATA II), and from October 2023 to February 2024 (DATA III) in order to systematically investigate the properties of $\psi(3770)$ and $D$ meson production and decays. Precisely determining the luminosity of this dataset is important for various purposes. It is essential for measuring the cross section of $e^{+}e^{-}\to\psi(3770)\to D\bar{D}$, calculating the normalization factors in strong-phase measurements of $D^{0}$ decays [1, 2], and reducing the systematic uncertainty of analyses using the single tag method [3]. Additionally, the luminosity is used to normalize the Monte Carlo (MC) sample size and to estimate continuous background in the $\psi(3686)$ dataset [4, 5]. This paper presents a measurement of the integrated luminosity in DATA I, DATA II, and DATA III using large angle Bhabha scattering events. This dataset is currently the world’s largest collection of $e^{+}e^{-}$ collision data at the $\psi(3770)$ resonance peak.
2 BESIII detector and Monte Carlo simulation
The BESIII detector [6] records energy-symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring [7] in the $E_{\rm cm}$ range from 2.0 to 4.95 GeV, with a peak luminosity of $1.1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $E_{\rm cm}=3.773\;\text{GeV}$ in 2023. BESIII has collected large data samples in this energy region [8, 9, 10]. The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1\leavevmode\nobreak\ {\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region is 60 ps [11].
Simulated data samples produced with a geant4-based [12] Monte Carlo (MC) package, which includes the geometric description of the BESIII detector and the detector response, are used to determine detection efficiencies and to estimate backgrounds. The QED processes are simulated with the Babayaga@NLO generator [13, 14, 15, 16]. The width of the $\psi(3770)$, beam energy spread, initial state radiation (ISR), and final state radiation (FSR) are considered in the simulation. The configuration parameters of the Bhabha signal MC ($e^{+}e^{-}\to(\gamma)e^{+}e^{-}$) are listed in Table LABEL:optionMC. As for other processes, the simulation models the beam energy spread and ISR in the $e^{+}e^{-}$ annihilations with the generator kkmc [17]. The inclusive MC sample is used to simulate all possible processes of $e^{+}e^{-}$ collision, including the production of $D\bar{D}$ pairs (including quantum coherence for the neutral $D$ channels), the non-$D\bar{D}$ decays of the $\psi(3770)$, the ISR production of the $J/\psi$ and $\psi(3686)$ states, and the continuum processes incorporated in kkmc [17].
All particle decays are modelled with evtgen [18] using branching fractions either taken from the Particle Data Group [19], when available, or otherwise estimated with lundcharm [20]. FSR from charged final state particles is incorporated using the photos package [21].
In the simulation above, the run-by-run based calibration is applied, taking into consideration the beam energy fluctuations and other time-dependent effects. The duration of each run is typically one hour. The calibration of $E_{\rm cm}$, which takes advantage of the small uncertainty in the known $D$ mass $m_{D}$ [19] and the excellent resolution of the MDC and EMC, is carried out using the following equations:
$E_{\rm cm}=2E_{D}\,,E^{2}_{D}=E_{0}^{2}+(m_{D}c^{2})^{2}-(M^{\rm fit}_{\rm BC}c^{2})^{2}.$ (2)
Here, $E_{0}$ is the uncalibrated beam energy, i.e., 3.773 GeV, and $M^{\rm fit}_{\rm BC}$ denotes the fitted peak of the beam-constrained mass of the $D$ mesons. This is calculated using $M^{\rm fit}_{\rm BC}c^{2}=\sqrt{E_{0}^{2}-p_{D}^{2}c^{2}}$, where $p_{D}$ is the momentum of the $D$ measured in the center-of-mass system of the $e^{+}e^{-}$ collision using the decays $D^{0}\to K^{-}\pi^{+}$, $D^{0}\to K^{-}\pi^{+}\pi^{+}\pi^{-}$ and $D^{+}\to K^{-}\pi^{+}\pi^{+}$. Essentially, Eq. 2 is equivalent to $E^{2}_{D}=p^{2}c^{2}+m^{2}_{D}c^{4}$. The distributions of $M_{\rm BC}$, instead of $p_{D}$, are utilized because $M_{\rm BC}$ offers better resolution and the fit to $M_{\rm BC}$ is more easily controlled. According to the above analysis, the calibrated $E_{\rm cm}$ varies by 2-3 MeV around the expected 3.773 GeV with an uncertainty of approximately $0.02$ MeV. Details of the fit and particle reconstruction are introduced in Ref. [22].
[…]
Configuration of the Babayaga@NLO generator used to simulate Bhabha events. Parameter Value Center-of-mass energy Calibrated $E_{\rm cm}$ Beam Energy Spread 0.97 MeV Minimum $\cos\theta$0.83 Maximum $\cos\theta$ 0.83
3 Method
The integrated luminosity of data is usually measured with three QED processes: $e^{+}e^{-}\to(\gamma)e^{+}e^{-}$, $e^{+}e^{-}\to(\gamma)\gamma\gamma$ and $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$. The symbol “($\gamma$)” represents the possible presence of photon(s) resulting from ISR or FSR. Since the uncertainties associated with reconstructing electrons are smaller than those of photons and muons, and the statistical uncertainty of Bhabha events are almost negligible, only Bhabha event is used to measure luminosity. Furthormore, Bhabha events at small angles, i.e., in the direction of the $e^{+}e^{-}$ beam, predominantly interact with the end cap region of the detector. This region exhibits more gaps and worse resolution compared to the barrel region. The generator simulation at small angles is also less precise. Given that large angle Bhabha events provide a negligible statistical uncertainty, we exclusively utilize them to measure the integrated luminosity of the data set. The large angle region is defined as $|\cos\theta|<0.83$, where $\theta$ is the polar angle of the final state electron (positron) relative to the beam direction.
The integrated luminosity of data is determined by
$\mathcal{L}=\sum_{i}\frac{N_{i}^{\rm obs}\times(1-\eta)}{\sigma_{i}\times\epsilon_{i}\times\epsilon^{\rm trig}}\,.$ (3)
In this equation, the index $i$ indicates different runs, and the integrated luminosity of the data set is obtained by summing over these runs. The variables $N_{i}^{\rm obs}$, $\sigma_{i}$, $\epsilon_{i}$, $\eta$, and $\epsilon^{\rm trig}$ represent the number of observed Bhabha-event candidates, the production cross section of the Bhabha process, the detection efficiency, the contamination rate, and the trigger efficiency for collecting events in the on-line data acquisition system, respectively. The Babayaga@NLO generator [13, 14, 15, 16] is employed, using the $E_{\rm cm}$ calibrated on a run-by-run basis, to calculate the cross section, generate signal MC events for the process $e^{+}e^{-}\to(\gamma)e^{+}e^{-}$, and estimate the detection efficiency.
4 Luminosity measurement
4.1 Event selection
Candidate Bhabha events are required to have exactly two oppositely charged tracks detected in the MDC. Each track is required to satisfy a distance of closest approach to the interaction point of less than 10 cm along the $z$ axis (the symmetry axis of the MDC) and less than 1 cm in the transverse plane. Additionally, the each candidate track must lie within the polar angle region $|\rm cos\theta|<0.8$ to ensure interaction with the barrel of the EMC.
To suppress the $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ background, the deposited energy $E_{\rm EMC}$ in the EMC by each track, shown in Fig. 4.1, must fall within the range $2.51.0<E_{\rm EMC}<$ 2.5 GeV. Furthermore, the sum of the momenta of the two tracks, shown in Fig. 4.1, is required to be larger than $0.9\times E_{\rm cm}$ to suppress background events, which mainly from processes involving a $J/\psi$ in the final state, e.g., $e^{+}e^{-}\to(\gamma)J/\psi$, $e^{+}e^{-}\to(\gamma)\psi(3686)\to(\gamma)J/\psi X$, and $e^{+}e^{-}\to\psi(3770)\to(\gamma)J/\psi X$. To eliminate the background from energetic cosmic rays, the momentum of each track must be less than $E_{\rm cm}/2+0.30$ GeV.
The two oppositely charged tracks in the candidate Bhabha scattering events are produced to be back-to-back. However, due to the presence of a magnetic field, their trajectories are bent, resulting in their respective shower clusters in the $xy$-plane of the EMC not being back-to-back. To quantify this angular difference, the variable $\delta\phi$ is defined as $|\phi_{1}-\phi_{2}|-180^{\circ}$, where $\phi_{1}$ and $\phi_{2}$ are the azimuthal angles of the two clusters in the EMC. The distribution of $\delta\phi$ is illustrated in Fig. 4.1 and a requirement of $5^{\circ}<|\delta\phi|<40^{\circ}$ is imposed. This requirement effectively eliminates the background from the $e^{+}e^{-}\to(\gamma)\gamma\gamma$ process.
[…]
The distribution of $E^{e^{+}}_{\rm EMC}$ versus $E^{e^{-}}_{\rm EMC}$ from a subset of the data sample. The cluster concentrated at the upper right corner is the Bhabha signal and the small cluster at the lower left corner is the $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ background. The horizontal and vertical bands are caused by ISR and FSR in Bhabha and $e^{+}e^{-}\to(\gamma)\mu^{+}\mu^{-}$ events.
[…]
The distribution of $p_{e^{+}}+p_{e^{-}}$ from a subset of the data sample. The background is not visible due to the very low contamination rate of $\eta=3\times 10^{-4}$. The pink arrow indicates the $>0.9\times E_{\rm cm}$ requirement.
[…]
The distribution of $\delta\phi$ between the selected $e^{+}$ and $e^{-}$ tracks from a subset of the data sample. The background is not visible due to the very low contamination rate of $\eta=3\times 10^{-4}$. The pink arrows indicate the $5^{\circ}<|\delta\phi|<40^{\circ}$ requirement.
4.2 Background estimation
The residual background events are from various sources, such as the ISR production of $J/\psi$ and $\psi(3686)$, $e^{+}e^{-}\to(\gamma)\gamma\gamma$, $e^{+}e^{-}\to\psi(3770)\to D\bar{D}$, $e^{+}e^{-}\to\psi(3770)\to$ non-$D\bar{D}$, and $e^{+}e^{-}\to$ continuum processes. These background contributions are estimated by analyzing the corresponding background MC events. The contamination rate for the candidate Bhabha events is calculated as $\eta=N^{\rm MC}_{\rm bkg}/(N^{\rm MC}_{\rm BB}+N^{\rm MC}_{\rm bkg})$, where $N^{\rm MC}_{\rm BB}$ and $N^{\rm MC}_{\rm bkg}$ are the numbers of Bhabha and background MC events that satisfy the selection criteria, respectively. These numbers are normalized according to individual cross sections. The resulting contamination rate is $\eta=3\times 10^{-4}$.
4.3 Detection efficiency and cross section
To estimate the detection efficiency for the Bhabha events, an $e^{+}e^{-}\to(\gamma)e^{+}e^{-}$ signal MC sample is generated for each run, using the calibrated $E_{\rm cm}$, with the Babayaga@NLO generator [13, 14, 15, 16]. The MC samples are generated within the range of $|\cos\theta|<0.83$, where $\theta$ represents the polar angle of the final state $e^{+}$ and $e^{-}$. By applying the same selection criteria as used in the data analysis to these MC samples, the efficiency is calculated as the ratio of the number of selected signal MC events to the generated number of events. The cross section for each run is also estimated using Babayaga@NLO, taking into account the calibrated $E_{\rm cm}$ of each run. A total of 2 billion signal MC events are generated, which corresponds to approximately the same size as the data. This large sample size allows us to neglect the uncertainty arising from the statistics of the signal MC sample. The detection efficiency and cross section within $|\cos\theta|<0.83$ at 3.773 GeV are determined to be 61.09% and 147.47 nb, respectively, with fluctuations on the order of $0.1\%$ for different runs due to variations in $E_{\rm cm}$.
4.4 Integrated luminosities
The number of observed Bhabha events for each run $N_{i}^{\rm obs}$ is determined by counting the events that satisfy the selection criteria outlined in Sec. 4.1. In total, $450.97\times 10^{6}$, $736.88\times 10^{6}$, and $379.45\times 10^{6}$ events are obtained for DATA I, DATA II, and DATA III, respectively. The trigger efficiency $\epsilon^{\rm trig}$ for collecting $e^{+}e^{-}\to(\gamma)e^{+}e^{-}$ events has been measured to be 100% with a statistical uncertainty of less than 0.1% [23].
Inserting the number of observed Bhabha events, the trigger efficiency, the contamination rate (Sec. 4.2), the detection efficiency, and the cross sections within $|\cos\theta|<0.83$ (Sec 4.3) into Eq. (3), the integrated luminosities are determined to be $4.995\pm 0.019$ fb-1 for DATA I, $8.157\pm 0.031$ fb-1 for DATA II, and $4.191\pm 0.016$ fb-1 for DATA III, where the statistical uncertainties are negligible and the systematic uncertainties will be discussed in the next subsection.
4.5 Systematic uncertainty
The systematic uncertainty arising from the MDC information, which includes the MDC tracking efficiency and the momentum requirement, is determined to be 0.29% by comparing the integrated luminosities measured with and without the MDC information. The distribution of $\delta\phi$ without the MDC information applied is shown in Fig. 4.5, and the efficiency increases by about 5%. The dominant background is $e^{+}e^{-}\to(\gamma)\gamma\gamma$, which is at the 0.15% level. The systematic uncertainty associated with the $E_{\rm EMC}$ requirements is estimated by varying the requirement from 1.0 to 0.9 or 1.1 GeV. This variation results in a change of 0.16% in the luminosity for both tracks. To estimate the systematic uncertainty caused by the $\cos\theta$ requirement, the integrated luminosity is determined with $|\cos\theta|<0.75$ or 0.70, and the difference from the standard selection of $|\cos\theta|<0.80$ is found to be 0.13% for both tracks, which is assigned as the corresponding systematic uncertainty. The uncertainty due to the $\delta\phi$ signal region selection is estimated to be 0.02% by comparing the integrated luminosities obtained by changing the lower limit from $5^{\circ}$ to $0^{\circ}$, or the higher limit from $40^{\circ}$ to $20^{\circ}$ or $30^{\circ}$. The uncertainty arising from the trigger efficiency [23] and the theoretically calculated cross section using the Babayaga@NLO generator [15, 24, 25] are both 0.1%. The uncertainties from the MC statistics and the contamination rate are negligible.
[…]
The distribution of $\delta\phi$ between the selected $e^{+}$ and $e^{-}$ tracks without MDC information from a subset of the data sample.
All contributions to the systematic uncertainty are summarized in Table LABEL:Tab:err_lum. The total systematic uncertainty is obtained by adding all individual contributions in quadrature.
5 Summary
By analyzing large angle Bhabha scattering events, we have measured the integrated luminosity of the $e^{+}e^{-}$ collision data collected at $E_{\rm cm}=3.773$ GeV from the years 2021 to 2024 with the BESIII detector at the BEPCII collider. These are crucial normalization factors for experimental studies of the production and decays of the $\psi(3770)$ and $D$ mesons. Using the cross sections of $\sigma(e^{+}e^{-}\to D^{0}\bar{D}^{0})=(3.615\pm 0.010_{\rm stat}\pm 0.038_{\rm syst})$ nb and $\sigma(e^{+}e^{-}\to D^{+}D^{-})=(2.830\pm 0.011_{\rm stat}\pm 0.026_{\rm syst})$ nb [26], one can obtain the numbers of $D^{0}\bar{D}^{0}$ and $D^{+}D^{-}$ pairs in data. All the numerical results are summarized in Table LABEL:tab:lum. Along with the $(2.932\pm 0.014)$ fb-1 data collected from 2010 to 2011 [27, 28], BESIII has accumulated data sets with a total integrated luminosity of $(20.275\pm 0.077)$ fb-1 at $E_{\rm cm}=3.773$ GeV. These results offer fundamental inputs for physics analyses based on these data samples [29, 30].
[…]
Numerical results for DATA I, DATA II, and DATA III along with data collected from 2010 to 2011 (DATA 2010) [27], the number of observed events ($N_{\rm obs}$), the integrated luminosity ($\mathcal{L}$), and the numbers of $D^{0}\bar{D}^{0}$ ($N_{D^{0}\bar{D}^{0}}$) and $D^{+}D^{-}$ ($N_{D^{+}D^{-}}$). Sample $N_{\rm obs}$ ($10^{6}$) $\mathcal{L}$ (fb-1) $N_{D^{0}\bar{D}^{0}}$ ($10^{6}$) $N_{D^{+}D^{-}}$ ($10^{6}$) DATA I 450.97 $4.995\pm 0.019$ $18.06\pm 0.21$ $14.14\pm 0.16$ DATA II 736.88 $8.157\pm 0.031$ $29.49\pm 0.34$ $23.08\pm 0.25$ DATA III 379.45 $4.191\pm 0.016$ $15.15\pm 0.18$ $11.86\pm 0.13$ DATA 2010 [27] 283.95 $2.932\pm 0.014$ $10.60\pm 0.13$ $8.30\pm 0.09$ TOTAL 1851.25 $20.275\pm 0.077$ $73.29\pm 0.84$ $57.38\pm 0.61$